a leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3 centimeters shorter than one and one-half times the length of the equal sides. the perimeter is 28.5 centimeters. Find the lengths of the sides.

Answer by checkley77(12844)

x+x+(1.5x-3)=28.5
2x+1.5x-3=28.5
3.5x=28.5+3
3.5x=31.5
x=31.5*1/3.5
x=9 ans. for the 2 equal sides.
1.5*9-3=13.5-3=10.5 ans. for the long side.
Proof:
9+9+10.5=28.5
28.5=28.5